It was claimed that 1 of 5 cardiologists take aspirin a day to prevent hardening of the arteries. Suppose that the claim is true. If 16 cardiologists are selected independently at random, let X be the number who take aspirin a day. (a) How is X distributed? (b) Give the values of E(x) and Var (x) (c) Determine P(X>5)
Answer #1

Solution:

a) X follow the binomial distribution with n=16 and probability of success p= 1/5= 0.2

b) E(X)=np= 16*0.2= 3.2 and

V(X)= np(1-p)= 16*0.2*0.8= 2.56

C) P(X > 5)= 1 - P(X ≤ 4) = 1 - 0.7982= 0.2018

Calculation:

The following information is provided: The population proportion of success is p = 0.2, also, 1-p=1 – 0.2 = 0.8, and the samp